How many years would it take for a debt of $10,915 to grow into $13,401 if the annual interest rate is 3.9% with daily compounding?
Answer #1

Ans 5.26 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
13401 / 10915 = (1 + 3.9%/365)^n
1.22775996335318 = 1.00010684931507^n
n = 1920 periods
or
n = 5.26 years